OpenStudy (anonymous):

Does the function satisfy the hypotheses of the Mean Value Theorem on the given interval? f(x) = x/x+6 [1, 12] If it satisfies the hypotheses, find all numbers c that satisfy the conclusion of the Mean Value Theorem. (Enter your answers as a comma-separated list. If it does not satisfy the hypotheses, enter DNE). Ive tried to do it multiple ways but i cant get the right answer. Please help!

jimthompson5910 (jim_thompson5910):

the function is \[\Large f(x) = \frac{x}{x+6}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

which x value makes the denominator zero?

OpenStudy (anonymous):

-6

jimthompson5910 (jim_thompson5910):

is -6 in the interval [1,12] ?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

so the function is continuous on [1,12] that allows us to use the MVT properly here

jimthompson5910 (jim_thompson5910):

if there was a discontinuity on [1,12] then we couldn't use the MVT

OpenStudy (anonymous):

okay, I understand that part, the equation part is whats given me a lot of trouble, i dont have examples like this one

jimthompson5910 (jim_thompson5910):

Mean Value Theorem If f(x) is continuous on [a,b] and differentiable on (a,b), then there exists at least one value of c such that \[\Large f \ '(c) = \frac{f(b) - f(a)}{b-a}\]

jimthompson5910 (jim_thompson5910):

what you have to do is find the secant slope through (1, f(1)) and (12,f(12)) then find f ' (x) set f ' (x) equal to the secant slope and solve for x

OpenStudy (anonymous):

so i would substitute 1 for a and 12 for b? that gives me 1?

jimthompson5910 (jim_thompson5910):

you said `that gives me 1?` what do you mean exactly? that's the secant slope you got?

OpenStudy (anonymous):

yeah if i substitute the a for 1 and 12 for b, that equals to 11/11 so 1? i dont think im doing this right

jimthompson5910 (jim_thompson5910):

that's not the correct secant slope

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{x}{x+6}\] \[\Large f(1) = \frac{1}{1+6}\] \[\Large f(1) = \frac{1}{7}\]

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{x}{x+6}\] \[\Large f(12) = \frac{12}{12+6}\] \[\Large f(12) = \frac{12}{18}\] \[\Large f(12) = \frac{2}{3}\]

OpenStudy (anonymous):

and then i would do the same for f(12)?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay I got that part!

jimthompson5910 (jim_thompson5910):

you should have this next \[\Large f \ '(c) = \frac{f(b) - f(a)}{b-a}\] \[\Large f \ '(c) = \frac{f(12) - f(1)}{12-1}\] \[\Large f \ '(c) = \frac{\frac{2}{3} - \frac{1}{7}}{12-1}\] \[\Large f \ '(c) = \frac{\frac{11}{21}}{11}\] \[\Large f \ '(c) = \frac{1}{21}\]

jimthompson5910 (jim_thompson5910):

now you need f ' (x)

OpenStudy (anonymous):

so then i would substitute 1/21 for x in this 6/(6+x)^2 right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you set them equal to one another and solve for x

OpenStudy (anonymous):

I set what equal to each other?

jimthompson5910 (jim_thompson5910):

1/21 and that f ' (x) you got

OpenStudy (anonymous):

i end up with 6=1/21x^2+4/7x+12/7

OpenStudy (anonymous):

what would i do w the other x?

jimthompson5910 (jim_thompson5910):

\[\Large f \ '(x) = \frac{6}{(x+6)^2}\] \[\Large \frac{1}{21} = \frac{6}{(x+6)^2}\] \[\Large 1*(x+6)^2 = 21*6\] \[\Large (x+6)^2 = 126\] keep going to solve for x

OpenStudy (anonymous):

got -6+3root14 and -6-3root 14?

jimthompson5910 (jim_thompson5910):

those are the correct solutions to the equation you now need to check if they lie in the interval [1,12]

OpenStudy (anonymous):

okay so it would only be -6+3root14

jimthompson5910 (jim_thompson5910):

yes since \(\large -6+3\sqrt{14} \approx 5.22\) which is in the interval [1,12]

jimthompson5910 (jim_thompson5910):

the other value is -17.22 which is not in that interval

OpenStudy (anonymous):

thank you so much!!